OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1)

OpenStudy (anonymous):

so, so first you want to solve for x in the first equation. can you do that?

OpenStudy (anonymous):

x=1+4y usin d 1st equation after substitutin x in equation 2 v get 2(1+4y)-9y=3,2+8y-9y=3,-y=1, y=-1 usin d value of y v can get x by pluggin it in any of d equations. x-4(-1)=1 x+4=1,x=3

OpenStudy (anonymous):

oops srry x is -3